OpenStudy (anonymous):

Given f(x) = 6x^2 - x - 12 and g(x) = 2x - 3, find the function (f/g)(x). Enter your answer as the following example: -10x+1.

OpenStudy (anonymous):

what does the (f/g)(x) mean

OpenStudy (anonymous):

this one you need to write a fraction first \[\frac{6x^2-x-12}{2x-3}\]

OpenStudy (anonymous):

so its dividing?

OpenStudy (anonymous):

\[f/g\] means \[\frac{f(x)}{g(x)}\] yes it is dividing

OpenStudy (anonymous):

yeah i got that equation but i dont know what to do from there..

OpenStudy (anonymous):

my guess is that you can actually divide, i.e. you can factor the numerator and cancel lets check it

OpenStudy (anonymous):

\[6x^2-x-12=(2x-3)(3x+4)\] so you get \[\frac{6x^2-2x-12}{2x-3}=\frac{(2x-3)(3x+4)}{2x-3}=3x+4\]

OpenStudy (anonymous):

the clue that you could factor and cancel came from this line here in your question "Enter your answer as the following example: -10x+1."

OpenStudy (anonymous):

oh, thanks again, lol i feel stupid.... :(

OpenStudy (anonymous):

since they did not expect you to put in a fraction, it was clear that the numerator had to factor as \((2x-3)(\text{something})\) otherwise there would be nothing to do

OpenStudy (anonymous):

don't feel stupid! this is just math, you learn it, you are not born with it

OpenStudy (anonymous):

haha yeah, im in a lesson on my online school and i dont understand the way they teach so i have to look it all up online. :( thanks for helping, i understand it a bit better than when they taught me..

OpenStudy (anonymous):

glad to help, keep posting on line classes are difficult, because there is no one to ask if you have a question

OpenStudy (anonymous):

wait so cactus high is on line?

OpenStudy (anonymous):

i might have more questions soon though